OpenStudy (anonymous):

A bicycle traveled 150 meters west from point A to point B. Then it took the same route and came back to point A. It took a total of 2 minutes for the bicycle to return to point A. What is the average speed and average velocity of the bicycle?

OpenStudy (anonymous):

S = 150*2=300meters t=2 min = 120 s average speed is whole distance divided by whole time: Vaverage =S/t=300/120=2.5 metters per second Velocity will be equal to 0 since it doesn't change it's coordination.